Find the unknown angle for the following right triangle:
Angle A Angle B Angle C
? 90 27

Answer by Alan3354(21599)

The sum of angles in a triangle is always 180.
180 = A+B+C
180 = A+90+27
180 = A+117
A = 63 degrees